1. What are the next three terms in the sequence?
–6, 5, 16, 27, … (1 point)
38, 49, 60
37, 47, 57
36, 45, 54
36, 46, 57
please check my answer I think it is c

Answer by Edwin McCravy(20054)

No it's not.  We notice that we add 11 each time to get the next number.
We subtract 2nd term minus 1st term, then 3rd term-2nd term, etc., and
observe the pattern.  Then follow the same pattern to form the next three. 

   sequence      2nd-1st    3rd-2nd    4th-3rd     pattern       next 3       
–6, 5, 16, 27,  5-(-6)=11   16-5=11    27-16=11  +11,+11,+11   38,49,60

Reason: 27+11=38, 38+11=49, 49+11=60

Edwin